# Invariance under Structure Translation as the Origin of Host Immune Capacity Conservation from Noether’s Theorem

### v2

@@ 5.5 Integrating the Components as Total Immune Capacity
The total Immune Capacity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ emerges from the integration of its components, representing a holistic measure of the host’s protective potential. It is this integrated quantity but not any single parameter in isolation that most accurately corresponds to the clinical notion of Immune Capacity. A high value of $\mathbf{I}$ can result from various combinations of its constituents: a large effective mass $m_{1}$, a high coupling efficiency $k$, a highly coordinated multi-organ response (represented by $Q$), or a robust innate immune response for the preemptive reduction of antigenic load, which attenuates the impulse $\Delta p_{2}$ presented to the adaptive system under an identical challenge. This formalism quantifies the system-level breadth and amplitude of the immune state, indicating that while acute responses to specific pathogens may vary, the global protective capacity is maintained for a given host condition.
@@ 6 The Measurement of Immune Capacity
The previous section established the relationship between the conserved quantity for immune recognition, $I_{\mathrm{Spatial}}$, the invariant for response intensity, $I_{\mathrm{Temporal}}$, and key parameters ($k,m_{1},q_{1},\dot{q}_{1},Q$). Based on this theoretical framework, the subsequent critical task is the practical measurement or inference of these parameters and the Immune Capacity $\mathbf{I}$.
@@ 6.1 Determination of Immune Parameters in Model Organisms
An experimental paradigm can be designed as follows: First, under homeostatic conditions (no antigenic challenge), $q_{2}$ and $\dot{q}_{2}$ are zero, and the generalized force $Q$ is zero, indicating a balance between immunological energy dissipation and input. The conserved quantity simplifies to:
$I_{\mathrm{Spatial}}=kp_{1}=km_{1}\dot{q}_{1}$ (29)
Consequently, the baseline $\dot{q}_{1}$ reflects the basal turnover rate of immune diversity, measuring it allows for an estimation of the fundamental Immune Capacity, $I_{\mathrm{Spatial}}$. For instance, parameters in a reference state (e.g., a healthy 42-day-old mouse) can be normalized ($k=1,m_{1}=1$) for initial quantification.
Subsequently, applying a standardized antigenic challenge (e.g., VLP-A in this study, with a defined reference impulse $p_{2}=1$) enables dynamic monitoring of the changes in $q_{1}$, $\dot{q}_{1}$, $\Delta p_{1}$, and the estimation of $I_{\mathrm{Temporal}}$ over time, thereby establishing a ”standard curve” of the systemic response.
@@ 6.2 Assessment of Immune Capacity in Healthy Human Populations
When assessing Immune Capacity $\mathbf{I}$ in healthy human populations, the constant presence of environmental antigens must be considered. Antigen exposure activates the immune system, leading to non-zero values of $q_{2}$ and $\dot{q}_{2}$ , which introduces an impulse $\Delta p_{2}$. This in turn influences $p_{1}$ and $\dot{q}_{1}$, ultimately causing a displacement in $q_{1}$. Upon reaching a steady state under persistent antigen exposure, $\dot{q}_{2}$, $p_{2}$ returns to zero, and the generalized force $Q=0$, indicating a balance between immune energy dissipation and input. Under these conditions, the conserved quantity still simplifies to:
$I_{\mathrm{Spatial}}=kp_{1}=km_{1}\dot{q}_{1}$ (30)
Thus, in the presence of continuous antigenic challenge, differences in immune steady states are primarily reflected in the immune diversity $q_{1}$, and Immune Capacity $\mathbf{I}$ still remains proportional to the baseline turn over rate of immune diversity, $\dot{q}_{1}$. Since administering standardized antigen challenges in human populations is impractical, direct measurement of parameters such as $k$, $m_{1}$, and $Q$ is challenging. A feasible alternative is to estimate the average $\dot{q}_{1}$ from large-scale population data, establishing baseline Immune Capacity $\mathbf{I}$ references stratified by factors including, but not limited to, geography, age, and ethnicity.
@@ 6.3 Immune Capacity Assessment in Diseased Populations
Different diseases correspond to distinct disease locations and pathological changes, generating different stimulation signals. Consequently, they produce distinct functional forms of $Q$, which ultimately act on the immune system and result in different immune states. Therefore, by comparing immune state differences with the healthy state, the disease course can be determined; and by comparing immune states across different diseases, the disease type can be identified. In our previous work, we estimated the velocity $\dot{q}_{1}$ from immune repertoire diversity shifts (displacement $S$), considering clonal expansion and new clone generation from somatic hypermutation as two primary displacement mechanisms. To quantify differences in immune states induced by different forms of $Q$, we employed the Wasserstein distance to measure the disparity between two immune distributions, using the minimal cost of distribution transformation as the immune distance. This approach successfully distinguished colorectal cancer from healthy controls, and Kawasaki disease from general fever in blinded data.
However, being based on curve-fitting and Markov chain modeling rather than first principles, the method remains phenomenological and mechanistically vague. In this article, we revisit the inference of immune capacity from a more systematic, top-down perspective.
@@ 6.4 Experimental Measurement of Key Immune Parameters
The theoretical framework establishes that the conserved quantity $I_{\mathrm{Spatial}}$ under homeostasis is related to the basal turnover rate of immune diversity, $dot{q}_{1}$, while $I_{\mathrm{Temporal}}$ quantifies the degree of deviation from homeostasis (i.e., the displacement in $q_{1}$). Practically, comprehensive enumeration of all immune cells is infeasible; parameters must be estimated from sampled proportions or frequencies. Thus, robust sampling and statistical inference methods are essential. This section outlines potential experimental approaches for measuring these parameters.
6.4.1 Measuring the Basal Turnover Rate of Immune Diversity $\dot{q}_{1}$
In vivo, cellular turnover rates can only be estimated as proportions or frequencies. An abnormally high value of basal turnover rates may indicate a low cell reserve, whereas a low value may suggest insufficient immune dynamics. Establishing a normal reference range requires large-scale population studies. The most common and practical approach involves single-time-point phenotypic analysis to infer cell renewal dynamics. Key indicators include:
Proportion and Absolute Count of Naive B and T Cells
: A healthy, competent immune system maintains a large naive B-cell and T-cell pool. A decline in their proportion or count often indicates age-related involution.
Proliferation and Differentiation Markers
: The proportion of progenitor cells at intermediate stages of differentiation from hematopoietic stem cells to naive cells can indicate generative potential.
Immune Repertoire Sequencing and Naive Cell Estimation
: High-throughput sequencing of BCR or TCR enables the estimation of the naive cell proportion based on sequence characteristics. Clonal sequences that are unexpanded (present as a single copy) and carry no somatic hypermutations are classified as naive. The ratio of such naive-origin cells to the total lymphocyte population provides a molecularly defined estimate of the naive cell pool, from which basal turnover rates can be inferred.
DNA Labeling and Tracking
: Pulse-labeling with BrdU or stable isotopes, followed by tracking the decay of the labeled cell fraction over time, allows direct calculation of a population’s half-life and turnover rate via mathematical modeling of the decay curve.
Telomere Length Analysis
: Telomeres shorten with each cell division. Thus, the average telomere length of a cell subset reflects its replicative history, with naive cells typically having longer telomeres.
@@ 6.4.2 Quantifying Deviation from Homeostasis via Immune Repertoire Sequencing
The degree of immune deviation is quantified by comparing the receptor repertoire of a sample against a defined homeostatic baseline. This is achieved through immune repertoire sequencing, a set of high-throughput techniques for profiling the TCR or BCR genes to determine clonal diversity, sequence identity, and frequency distributions. Established methods include multiplex PCR, 5’-RACE (Rapid Amplification of cDNA Ends), and single-cell V(D)J sequencing, each offering distinct advantages in capturing repertoire breadth or paired-chain information.
In our previous research [Citation X], we employed a 5’-RACE-based protocol incorporating a template-switch mechanism. This method was selected for its efficiency in capturing the full-length V(D)J transcriptome from limited RNA input while minimizing amplification bias, thereby providing a quantitative and representative profile of the clonal landscape for robust comparative analysis.
@@ 6.5 Sample Size Considerations for Immune Parameter Measurement
Conventional flow cytometry (assaying millions of cells) suffices to estimate cell proportions for the adaptive immune system reliably. To stably estimate the distribution shape and compute the difference between distributions of highly diverse immune repertoire ($10^{20}$ unique sequences), which follows a scale-free power-law distribution[], futher mathematical tools are needed, for conventional diversity indices are unsuitable for power-law distributions of immune repertoire[Ref]. Our previous research utilized the Wasserstein distance to compare repertoires, which quantifies the minimal ”cost” to transform one distribution into another. This metric is well-suited for power-law distributions. We demonstrated that approximately $10^{4}$ high-quality sequences are sufficient to stably fit the power-law distribution and enable accurate distance calculation, a finding validated in blinded studies discriminating Kawasaki disease from colorectal cancer.
@@ 7 Microscopic Stochasticity and Macroscopic Immune Patterns
The existence of immunological memory implies that upon returning to homeostasis after an immune response, the potential energy surface of the system to that specific antigen is permanently altered compared to its naive state, manifesting as a reduced average energy barrier against specific antigens. The physical basis of this optimization is somatic hypermutation of antibody genes—an inherently stochastic process. The affinity maturation of a B cell clone can be modeled as a random walk within a one-dimensional affinity landscape, where clones with decreased affinity are selected against (cell death), and those with increased affinity are positively selected for expansion. The magnitude of affinity change resulting from each single mutation event follows a Gaussian distribution $N(\mu,\sigma)$ where minute changes are highly probable, while large-affinity leaps or drops are rare. In our model, this stochastic process corresponds to gradually optimizing the antibody structure to refine the system’s evolutionary trajectory, formally adjusting variable $q_{3}$ in system dynamics. Critically, although the specific antibody sequence for any single clone is stochastic, the overall efficiency of the affinity maturation process and the final, macroscopic level of protective immunity achieved are constrained by the system’s key parameters ($k$, $m_{1}$) and the conserved quantity $\mathbf{I}$. Thus, idefines the boundary and potential of the system’s statistical behavior.
@@ 8 Conclusion and Discussion
Our work does not seek to replace existing clinical measures but to explain and unify them under a single theoretical roof. The conserved quantity II provides a physical interpretation for why a ’strong’ or ’weak’ immunity manifests in certain ways. It explains clinical observations as natural consequences of conservation laws such as why a robust system responds calmly to a challenge while a frail system overreact.
@@
The quantity $\mathbf{I}$ have two components, $(I_{\mathrm{Spatial}}$ and $I_{\mathrm{Temporal}})$, they formally encapsulate the system’s protective breadth and response intensity respectively. $(I_{\mathrm{Spatial}}$ possesses the physical dimension of action $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$, $I_{\mathrm{Temporal}})$ possesses the dimension of energy $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$.

### v3

@@ 5.5 Integrating the Components as Total Immune Capacity
The total Immune Capacity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ emerges from the integration of its components, representing a holistic measure of the host’s protective potential. It is this integrated quantity but not any single parameter in isolation that most accurately corresponds to the clinical notion of Immune Capacity. A suitable value of $\mathbf{I}$ can result from various combinations of its constituents: a large total cell number $N$, a complete immune communication network $\eta$, an appropriate threshold for immune activation signals $\mu$, a highly coordinated multi-organ response (represented by $Q$), or a robust innate immune response for the preemptive reduction of antigenic load, which attenuates the impulse $\Delta p_{2}$ presented to the adaptive system under an identical challenge. This formalism quantifies the system-level sensitivity and amplitude of the immune state, indicating that while acute responses to specific pathogens may vary, the global protective capacity is maintained for a given host condition.
@@ 6 The Measurement of Immune Capacity
The previous section established the relationship between the conserved quantity for immune recognition, $I_{\mathrm{Spatial}}$, the invariant for response intensity, $I_{\mathrm{Temporal}}$, and key parameters ($k,N,\eta,\mu,\dot{q}_{1},Q$). Based on this theoretical framework, the subsequent critical task is the practical measurement or inference of these parameters and the Immune Capacity $\mathbf{I}$.
@@ 6.1 Determination of Immune Parameters in Model Organisms
An experimental paradigm can be designed as follows: First, under homeostatic conditions (no antigenic challenge), $\dot{q}_{2}$ is zero, and the generalized force $Q$ is zero, indicating a balance between immunological energy dissipation and input. The conserved quantity simplifies to:
$I_{\mathrm{Spatial}}=k_{2}p_{1}=\frac{K_{2}N\mu\dot{q_{1}}}{\eta}$ (30)
Measuring the basal turnover rate of immune diversity $\dot{q}_{1}$, the existing total immune cell number $N$, the intrinsic activation threshold and signaling efficiency $\mu$ and the structural integrity of the immune network $\eta$, allows for an estimation of the fundamental Immune Capacity, $I_{\mathrm{Spatial}}$. For instance, parameters in a reference immune state (e.g., a healthy 42-day-old mouse) can be normalized ($N=1,\mu=1,\dot{q_{1}}=1,\eta=1$) for initial quantification.
Subsequently, applying a standardized antigenic challenge (e.g., VLP-A in this study), with a defined immunogenicity $G$, a standard $K_{1}=1,K_{2}=1$, and a impulse $p_{2}=1$ enables dynamic monitoring of the changes in $\dot{q}_{1}$, $\Delta p_{1}$, and the estimation of $I_{\mathrm{Temporal}}$ over time, thereby establishing a standard baseline of the systemic response.
@@ 6.2 Assessment of Immune Capacity in Healthy Human Populations
When assessing Immune Capacity $\mathbf{I}$ in healthy human populations, the constant presence of environmental antigens must be considered. Continuous antigen exposure activates the immune system, resulting in non-zero values of the antigenic stimulation $q_{2}$ and $\dot{q}_{2}$ , which introduces a small yet persistent generalized momentum $p_{2}$. This in turn influences $p_{1}$ and $\dot{q}_{1}$, ultimately leading to a sustained displacement in $q_{1}$. Under conditions of persistent antigenic pressure, the system eventually reaches a dynamic equilibrium where the net generalized force returns to zero ($Q=0$), signifying a balance between energy dissipation within the immune system and energy input from antigenic stimulation. Consequently, the immune system resides in a state that is slightly yet stably displaced from a theoretical, antigen-free homeostasis.
Given the practical and ethical constraints of administering standardized antigen challenges in human studies, direct measurement of parameters such as $K_{1}$, $K_{2}$, $N$. $\mu$, $\eta$, $p_{2}$ and $Q$ is challenging. A viable alternative is to leverage high‑throughput omics profiling combined with large‑scale population datasets to indirectly infer the ensemble averages of these parameters. This approach enables the establishment of stratified baseline references for Immune Capacity $\mathbf{I}$, accounting for influential factors including geography, age, and ethnicity.
@@ 6.3 Immune Capacity Assessment in Diseased Populations
Different diseases correspond to distinct disease locations and pathological changes with different immunogenicity $G$, generating different stimulation signals. Consequently, they produce distinct functional forms of $Q$, which ultimately act on the immune system and result in different immune states. Therefore, by comparing immune state differences with the healthy state, the disease course can be determined; and by comparing immune states across different diseases, the disease type can be identified.
@@ 6.4 Experimental Measurement of Key Immune Parameters
The theoretical framework establishes that the conserved quantity $I_{\mathrm{Spatial}}$ under homeostasis quantifies the difficulty of altering the immune system’s dynamical state, while $I_{\mathrm{Temporal}}$ quantifies the degree of deviation from homeostasis (i.e., the displacement in $q_{1}$ and the velocity $\dot{q_{1}}$). In practice, comprehensive enumeration of all immune cells is infeasible; therefore, parameters must be estimated from sampled proportions or frequencies. Thus, robust sampling and statistical inference methods are essential. This section outlines potential experimental approaches for measuring these parameters.
6.4.1 Measuring the sensitivity of immune recognition
The quantity $I_{\mathrm{Spatial}}$ governs the sensitivity of the immune response to antigenic challenges and depends on several key parameters: the immunogenicity $G$, the strength of antigenic stimulation $p_{2}$, the total number of immune cells $N$, the per-cell activation barrier $\mu$, the functional and structural coherence of the immune signaling network $\eta$, and the cellular turnover rates $\dot{q_{1}}$.
Experimental estimation of these parameters requires a combination of quantitative assays and mathematical modeling. Immunogenicity $G$ can be assessed through in vitro stimulation assays that measure the dose-response relationship between antigen exposure and immune-cell activation (e.g., proliferation, cytokine secretion). The antigenic impulse $p_{2}$ may be approximated by the magnitude of early immune-cell expansion or antibody titers following controlled antigen exposure. Total immune-cell numbers $N$ are routinely obtained via flow-cytometric immunophenotyping of peripheral blood or lymphoid tissues, coupled with total lymphocyte counts. The activation barrier $\mu$ correlates with the mechanical force threshold for cell triggering, as demonstrated in studies where memory B cells exhibited ultra-low force thresholds due to PIP2-enriched IgG-BCR microdomains[]. The immune-network integrity $\eta$ can be estimated by quantifying the coordination among different immune cell subsets upon stimulation, for example, through the correlation of cytokine secretion profiles across cell types or the synchronicity of activation markers in response to antigen challenge. Finally, the rate of antigenic change $\dot{q_{2}}$ is context-dependent; in acute infections it may be estimated from pathogen replication kinetics (e.g., viral load doubling time), whereas in chronic settings it could correspond to the rate of antigenic drift. Integrating these measurements within a dynamical model allows the composite sensitivity $I_{\mathrm{Spatial}}$ to be estimated, providing a quantitative basis for comparing immune responsiveness across individuals or conditions.
@@ 6.4.2 Quantifying Deviation from Homeostasis via Immune Repertoire Sequencing
The extent of immune deviation is measured by contrasting a sample’s receptor repertoire against a reference, which may be another sample or a predetermined homeostatic baseline. Our laboratory previously introduced a computational strategy to define a distance metric between two distinct immune repertoire states. This immunological distance captures both its directional bias (for instance, whether the repertoire shifts toward clonally restricted populations or maintains a balanced, polyclonal composition) and its magnitude (which can be gauged by indicators such as the fraction of germinal-center B cells, the prevalence of activated T cell subsets, and the accrual of somatic hypermutations).
We mathematically framed this distance as the minimum ”cost” required to evolve one immune configuration into another. Conceptually, transforming a state characterized by limited clonality and a low mutational burden into one featuring broad diversity and a higher mutation load entails reshaping the initial clone distribution toward the target profile, while factoring out common elements. This cost is effectively computed using an optimal transport metric, specifically the Wasserstein distance, which quantifies the minimal effort to reconfigure one probability distribution into a different one.
A central consideration is that the biological expenditure for generating clones with varying attributes is not uniform. For example, producing a B-cell lineage carrying three somatic mutations incurs a greater cost than generating one with a single mutation. To account for this, we implemented an in silico, Markov chain-based simulation of repertoire evolution. This model provides empirical estimates for the transition costs between disparate cellular states. The final formulation merges principles from energy landscape theory with optimal transport theory, which portrays the relative ease or difficulty of state transitions and identifies the most efficient evolutionary trajectory. This synthesis bridges abstract computational modeling with tangible immune biology. The framework’s validity was demonstrated through its ability to discriminate mice across varying health strata and to clinically differentiate pediatric febrile illness from Kawasaki disease in a blinded, training-free manner, underscoring its robust predictive capacity and general applicability.
@@ 6.5 Sample Size Considerations for Immune Parameter Measurement
Conventional flow cytometry (assaying millions of cells) suffices to estimate cell proportions for the adaptive immune system reliably. To stably estimate the distribution shape and compute the difference between distributions of highly diverse immune repertoire ($10^{20}$ unique sequences), which follows a scale-free power-law distribution[], futher mathematical tools are needed, for conventional diversity indices are unsuitable for power-law distributions of immune repertoire[Ref]. Our previous research utilized the Wasserstein distance to compare repertoires, which quantifies the minimal ”cost” to transform one distribution into another. This metric is well-suited for power-law distributions. We demonstrated that approximately $10^{4}$ high-quality sequences are sufficient to stably fit the power-law distribution and enable accurate distance calculation.
@@ 7 Microscopic Stochasticity and Macroscopic Immune Patterns
The existence of immunological memory implies that upon returning to homeostasis after an immune response, the potential energy surface of the system to that specific antigen is permanently altered compared to its naive state, manifesting as a reduced average energy barrier against specific antigens. The physical basis of this optimization is somatic hypermutation of antibody genes an inherently stochastic process. The affinity maturation of a B cell clone can be modeled as a random walk within a one-dimensional affinity landscape, where clones with decreased affinity are selected against (cell death), and those with increased affinity are positively selected for expansion. The magnitude of affinity change resulting from each single mutation event follows a Gaussian distribution $N(\mu,\sigma)$ where minute changes are highly probable, while large-affinity leaps or drops are rare. In our model, this stochastic process corresponds to gradually optimizing the antibody structure to refine the system’s evolutionary trajectory, formally adjusting variable $q_{3}$ in system dynamics. Critically, although the specific antibody sequence for any single clone is stochastic, the overall efficiency of the affinity maturation process and the final, macroscopic level of protective immunity achieved are constrained by the system’s key parameters and the conserved quantity $\mathbf{I}$. Thus, idefines the boundary and potential of the system’s statistical behavior.
@@ 8 Conclusion and Discussion
The present framework models the immune system as operating in a time‑averaged, steady state and does not explicitly incorporate dynamic regulatory factors such as circadian rhythms or hormonal cycles. These factors are known to modulate immune reactivity on timescales ranging from hours to months and could be formally incorporated as time-varying components of the generalized forces Qacting on the system. By adopting a mean field, steady state representation, we effectively average over such higher frequency or periodic modulations, thereby focusing on the underlying structural and parametric determinants of immune capacity that persist across temporal fluctuations. This simplification is justified when studying coarse-grained, population level differences in immune competence, but future refinements of the model could explicitly introduce oscillatory or episodic forcing terms to capture the dynamic tuning of immune responsiveness by endocrine and neuro immune circuits.
@@
Our work does not seek to replace existing clinical measures but to explain and unify them under a single theoretical roof. The conserved quantity $I$ provides a physical interpretation for why a ’strong’ or ’weak’ immunity manifests in certain ways. It explains clinical observations as natural consequences of conservation laws such as why a robust system responds calmly to a challenge while a frail system overreact.
@@
The quantity $\mathbf{I}$ have two components, $I_{\mathrm{Spatial}}$ and $I_{\mathrm{Temporal}}$, they formally encapsulate the system’s protective sensitivity and response intensity respectively. $(I_{\mathrm{Spatial}}$ possesses the physical dimension of action $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$, $I_{\mathrm{Temporal}})$ possesses the dimension of energy $\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$.